OpenStudy (anonymous):

Write the equation y - 2 = 4(x + 5) in standard form. A. 4x + y = 7 B. 4x - y = 3 C. 4x - y = -22

OpenStudy (anonymous):

C

OpenStudy (anonymous):

y-2=4x+20 y=4x+22 y-4x=22 4x-y=-22 So your answer is C